OpenStudy (anonymous):

A contractor employed 210 men to build a house in 60 days. After 12 days,he was joined by 70 more men. In how many days will the remaining work will be finished ?

OpenStudy (anonymous):

No. worker is inversely proportional to the no. day. Therefore, 280/210=60/x 280x=210*60 x=12600/280 x=45 days The remaining work will be finished in 45 days.